# A Bias-Free Training Paradigm for More General AI-generated Image Detection

### v1

@@ 3.2 Metrics
4 Proposed Method
To realize and test our bias-free training paradigm we:
  • build a dataset consisting of real and generated fake images, where the latter are well aligned with their real counterparts but include the forensic artifacts of the diffusion-based generation process. The dataset is then enriched through several forms of augmentation, including locally inpainted images, and comprises eventually 309,102 generated images.
  • use this aligned dataset to fine-tune a Vision Transformer (ViT)-based model end-to-end. The model processes the entire ($512\times 512$) input image, so we avoid resizing and cropping operations altogether during fine-tuning. At inference time, $512\times 512$ crops are extracted from the image (if the image is larger we average the results of multiple crops).
Specifically, we use the image embedding network from the SigLIP model proposed in [49]. This model is pretrained for image-text matching but employs a pairwise sigmoid loss instead of the softmax loss function used in CLIP models. This modification improves memory efficiency, enabling the use of larger batch sizes and input dimensions. The dataset instead is created starting from the images collected from the training set of MS-COCO dataset [26], for a total of 51,517 real images.
To ensure fake images semantically match the content of real images, we exploit the conditioning mechanism of Stable Diffusion models that allows us to control the synthesis process through a side input, which can be a class-label, a text or another image. The side input is firstly projected to an intermediate representation by a domain specific encoder, and then feeds the intermediate layers of the autoencoders for denoising in the embedding space. After several denoising steps, a decoder is used to obtain the conditioned synthetic image from embedded vector (See Fig. 1). In our self-conditioned generation, we use the inpainting diffusion model of Stable Diffusion 2.1 [36], that has three side inputs: the reference image, a binary mask of the area to inpaint, and a textual description. Using an empty mask, we induce the diffusion steps to regenerate the input, that is, to generate a new image with exactly the same content of the input image. For the content augmentation process, we use the Stable Diffusion 2.1 inpainting method to replace an object with a new one, chosen from the same category or from a different one. Moreover, as shown in Fig. 4, besides the default inpainting, which regenerates the whole image, we consider also a version where the original background is restored. Note that during training, we balance the real and fake class taking an equal number of images from each.
In the following, we present our ablation study. To avoid dataset bias we use the WildRF [6] and Synthbuster [2] datasets, as explained in Section 3. In addition, we test on 1000 FLUX and 1000 Stable Diffusion 3.5 images, to verify performance on some of the latest synthetic generators.
@@ 4.1 Influence of content alignment
In Tab. 2 we show the performance achieved with different dataset alignment strategies, as described in Fig. 3. Note that all variants are trained with standard augmentations, including blurring and JPEG compression, as proposed in [43]. From the Table, we can observe that there is a large gain in terms of both AUC ($\simeq$10%) and balanced accuracy ($\simeq$20%) when upgrading from a text-driven generation (first row) to a solution where real and fake images share the semantic content, both reconstructed and self conditioned (see Fig. 3 for approaches). The proposed solution of using a diffusion pass demonstrates further improvement across all the evaluation metrics. This is highlighted in Fig. 5, where we show the power spectra evaluated by averaging the difference between the real and reconstructed images and the real and self-conditioned images. We observe that self-conditioned generation introduces forensic artifacts even at the lowest frequencies, indicating a detector trained on such images can exploit inconsistencies on a broader range of frequencies.
Figure 6: Robustness analysis in terms of balanced Accuracy carried out on nine generators of Synthbuster dataset [2].
Figure 7: Balanced accuracy of B-Free under various augmentation strategies evaluated from the initial online post. We analyzed multiple web-scraped versions of real and fake images that went viral on internet.
@@ 4.2 Effect of content augmentation
Overall, it is evident from Tab. 2 that augmentation plays a critical role in enhancing model generalization and this can be appreciated especially by looking at balanced accuracy and calibration measures. In fact, all of them improve significantly, while AUC is almost constant. More specifically, accuracy increases on average from around 83% to 89% thanks to our content-based augmentation. The most significant gains are observed on DALL·E 2, DALL·E 3 and FLUX that, probably, differ the most from Stable 2.1 in terms of architecture and hence require a stronger augmentation strategy to generalize.
@@
Ours SigLip 512 e2e 95.4 89.9 0.28 .066
Ours SigLip 512 LP 63.7 52.2 0.91 .285
Ours CLIP/ViT-L 224 e2e 75.2 73.9 0.66 .225
LDM [10] CLIP/ViT-L 224 e2e 63.9 48.5 3.39 .487
ProGAN [43] CLIP/ViT-L 224 e2e 54.7 45.2 4.85 .525
Table 3: Ablation study on the influence of the network architecture and the training data. We compare our solution, SigLip model trained end-to-end with content augumentation, with: fine-tuning (FT) of the last linear layer (linear probing, LP) a CLIP-based network and two well-known public training datasets: one based on ProGAN [43] and the other on Latent Diffusion [10].
@@
Table 4: Comparison with SoTA methods in terms of balanced Accuracy and balanced NLL across different generators. Note that AntifakePrompt [8] provides only hard binary labels hence calibration measures cannot be computed.
Robustness. In Fig. 6, we analyze the impact of our content augmentation, assessing robustness under various operations: JPEG compression, resizing, and blurring. Compared to using only self-conditioned images, all three proposed variants of augmentation offer a clear advantage, especially when resizing is applied. However, we can notice that in this experiment, the strongest augmentation strategy (inpainted++) may appear useless. Indeed, the relevance of such aggressive augmentation is more evident when analyzing images over the web. In fact, distinguishing real from synthetic images on social networks may be especially challenging due to the presence of multiple re-posting that impair image quality over time. A recent study conducted in [20] analyzed the detector behavior on different instances of an image shared online, showing that the performance degrades noticeably in time due to repeated re-posting. To better understand the impact of our augmentation strategies on such images, we collected a total of 900 real/fake images that went viral, including several versions of the same real or fake image. In Fig. 7 we show the balanced accuracy for our augmentation strategies, evaluated in different time periods from the first online post. In all cases the balanced accuracy gets significantly worse after just one day, but our strategy with more aggressive augmentation succeeds in stabilizing the performance to a very good level, always above $85\%$. This analysis makes clear that when testing the detector on images from the web a more intense augmentation inpainted++ is needed, otherwise the variant inpainted+ is sufficient. For remaining experiments, Ours refers to the detector trained using inpainted+ augmentation.
4.3 Influence of training data
@@
We conduct additional experiments to gain deeper insights into the impact of the chosen architecture and the proposed training data on the same datasets shown in Tab. 2. First, we compare our adopted model, SigLIP trained end-to-end, with an alternative fine-tuning strategy that involves training only the final linear layer, known as linear probing (LP) that is largely adopted in the literature [31, 11]. From Tab. 3 we can see that this latter solution does not perform well. One possible explanation is that features from last layer capture high-level semantics, while our dataset is built to exploit low-level artifacts that derive from first and intermediate layers [10, 21]. Then we use a different architecture, a CLIP-based model that works with an input size equal to $224\times 224$, and vary the training dataset by including two well known datasets largely used in the literature, one based on ProGAN [43] and the other on Latent Diffusion [10]. First we note that our training paradigm achieves the best performance over all the metrics, with a very large gain ($+17\%$ in bAcc). This is due to the adoption of a model that can handle a larger input size up to $512\times 512$ and hence can process at once the entire image of our training set.
@@
Table 5: AI-generated image detection methods used for comparison and whose code is made publicly available. We specify the source and size of the training dataset, and whether augmentation is applied.
Figure 8: Average performance in term of AUC and balanced accuracy of SoTA methods on four datasets: Synthbuster, GenImage, FakeInversion, SynthWildX.
5 Comparison with the State-of-The-Art
In this Section, we conduct a comparison with SoTA methods on 27 diverse synthetic generation models. To ensure fairness, we include only SoTA methods with publicly available code and/or pre-trained models. The selected methods are listed in Table 5 and are further described in the supplementary material together with additional experiments.
A first experiment is summarized in Tab. 4 with results given in terms of balanced accuracy and NLL. Most of the methods struggle to achieve a good accuracy, especially on more recent generators. Instead, B-Free obtains a uniformly good performance on all generators, irrespective of the image origin, whether they are saved in raw format or downloaded from social networks, outperforming the second best (see last column) by $+14.3\%$ in bACC and $+58\%$ in NLL. Then, we evaluate again all methods on GenImage (unbiased), FakeInversion [7], and SynthWildX [11]. As these datasets encompass multiple generators, we only report the average performance in Tab. 6. On these additional datasets, most methods provide unsatisfactory results, especially in the most challenging scenario represented by SynthWildX, with images that are shared over the web. The proposed method performs well on all datasets, just a bit worse on FakeInversion. Finally in Fig. 8 we study how AUC compares with balanced accuracy for all the methods over several datasets. We observe that some methods, like NPR and LGrad, present a clear non-uniform behavior, with very good performance on a single dataset and much worse on the others. This seems to suggest that these methods may not be truly detecting forensic artifacts, instead are rather exploiting intrinsic biases within the dataset. Differently, the proposed method presents a uniform performance across all datasets and a very small loss between AUC and accuracy.
@@ 7 Conclusions
In this work, we propose a new training paradigm for AI-generated image detection. First of all, we empirically demonstrate the importance of pairing real and fake images by constraining them to have the same semantic content. This helps to better extract common artifacts shared across diverse synthetic generators. Then we find that using aggressive data augmentation, in the form of partial manipulations, further boosts performance both in term of accuracy and of calibration metrics. This is extremely relevant especially when working in realistic scenarios, such as image sharing over social networks. Our findings emphasize that careful dataset curation and proper training strategy can be more impactful compared to developing more complex algorithms. We hope this work will inspire other researchers in the forensic community to pursue a similar direction, fostering advancements in bias-free training strategies.
@@ Acknowledgments.
In this supplementary document, we report more details about our implementation (Sec. 8). Moreover, we briefly describe the state of the art methods we compare to (Sec. 9), and give more details about the calibration metrics used in the experiments (Sec. 10). We also provide additional ablation results in Sec. 11 and carry out further experiments on generalization (Sec. 12) on additional publicly available datasets. Furthermore, we provide more results on the robustness of our approach compared with SoTA methods (Sec. 13).
8 Implementation Details
Training strategy. The proposed model leverages the SigLIP 512$\times$512 image embedding network as its backbone, followed by two fully connected layers. The model is trained end-to-end using the binary cross-entropy loss function on an NVIDIA A100 GPU. The training process employs the ADAM optimizer with a learning rate of 1e-6, a weight decay of 1e-6, and a batch size of 24. During training, the balanced accuracy is evaluated on a validation set every 3435 iterations. Early stopping is applied to prevent overfitting: training is completed if the validation balanced accuracy does not improve by at least 0.1% over five consecutive evaluations.
Test strategy. If the test image is less than 512 pixels, padding is applied after patch embedding. Otherwise, we average the logit score over multiple crops to analyze the whole image.
@@ 8 Implementation Details
Training Dataset. Here we give more deatils on how we built our dataset. Starting from the MS-COCO training set, consisting of 118K images with 80 categories of objects, we first discarded images with licenses different than Creative Commons. Before editing the images, we extracted the largest central crop, which allows us to retain most of the semantic content of the original image. We discarded images where objects are not present and ended up with a pristine source of 51,517 images. For content augmentation, we replaced the selected object with an object generated from the same category using the COCO segmentation mask, and from a different category using a rectangular box. We took care to not affect too much the realism of the content, so for the “different category” case the object is changed with one from a similar category, that belongs to the same COCO supercategory. In this scenario, the only exception is the category person, which does not have a supercategory and it is therefore replaced with a random object. As mentioned in the main paper, besides the default inpainting, we also consider a version where we take the pixels of the object from the generated image, and the pixels of the background from the original one. We did the same with the self-conditioned image, restoring the background with original pixels. Therefore, we ended up with six fake versions for each real image (Fig. 4 of the main paper).
9 SoTA Methods
Below we provide a brief description of the methods we included in the comparison in Section 5 of our main paper. The training datasets used by these methods are indicated in Table 5 of the main paper.
@@ 9 SoTA Methods
Ours 99.8 / 96.4 99.7 / 96.4 92.9 / 83.9 99.0 / 95.2 99.9 / 96.9 94.8 / 87.7 93.4 / 86.8 95.7 / 90.3 82.3 / 75.5 96.0 / 90.3 95.4 / 89.9
[…]
NLL / ECE RINE ProGAN [43] 5.03 / .410 1.99 / .236 0.77 / .135 20.5 / .546 0.36 / .053 10.1 / .506 1.22 / .149 6.51 / .454 2.46 / .242 5.23 / .393 5.43 / .312
LDM [10] 0.26 / .094 0.18 / .111 0.35 / .132 1.44 / .350 0.52 / .115 0.51 / .103 0.33 / .126 0.61 / .107 0.51 / .065 0.55 / .115 0.53 / .132
Ours 0.23 / .141 0.22 / .147 0.31 / .090 0.46 / .055 0.30 / .092 0.38 / .059 0.42 / .053 0.50 / .073 0.56 / .134 0.47 / .045 0.39 / .089
[…]
Ours 0.09 / .039 0.10 / .036 0.44 / .099 0.13 / .023 0.08 / .042 0.34 / .065 0.40 / .077 0.28 / .067 0.70 / .155 0.27 / .054 0.28 / .066
Table 7: Ablation study on the influence of the training data. We compare the performance of a SoTA method, RINE [21], by varying the training data: ProGAN, LDM and our dataset. Then we train our model, SigLip [49], on a publicly available dataset, D3 [3], that includes 4 generators from the Stable Diffusion family. Performance are presented in terms of AUC/Accuracy (top) and ECE/NLL (bottom).
@@
AntifakePrompt [8]. It relies on a visual question-answering (VQA) tool, InstructBLIP. The VQA is used with a fixed question, ”Is this photo real?”, and fine-tuned to provide accurate responses (”Yes” or ”No”) using a soft prompt tuning technique. Note that the method provides hard binary predictions hence only balanced Accuracy can be computed.
@@
CoDE [3]. CoDE trains a Vision Transformer using a contrastive loss similar to CLIP. However, while CLIP aims to learn features for text-image matching, CoDE aims at obtaining an embedding space where real and fake images are effectively separated.We report results using CoDE in combination with a the linear classifier.
10 Calibration Metrics

### v2

@@ 4 Proposed Method
To ensure fake images semantically match the content of real images, we exploit the conditioning mechanism of Stable Diffusion models that allows us to control the synthesis process through a side input, which can be a class-label, a text or another image. The side input is firstly projected to an intermediate representation by a domain specific encoder, and then feeds the intermediate layers of the autoencoders for denoising in the embedding space. After several denoising steps, a decoder is used to obtain the conditioned synthetic image from embedded vector (See Fig. 1). In our self-conditioned generation, we use the inpainting diffusion model of Stable Diffusion 2.1 [39], that has three side inputs: the reference image, a binary mask of the area to inpaint, and a textual description. Using an empty mask, we induce the diffusion steps to regenerate the input, that is, to generate a new image with exactly the same content as the input image. For the content augmentation process, we use the Stable Diffusion 2.1 inpainting method to replace an object with a new one, chosen from the same category or from a different one. Moreover, as shown in Fig. 4, besides the default inpainting, which regenerates the whole image, we consider also a version where the original background is restored. Note that during training, we balance the real and fake class taking an equal number of images from each.
In the following, we present our ablation study. To avoid dataset bias we use WildRF and Synthbuster. In addition, we test on 1000 FLUX and 1000 Stable Diffusion 3.5 images, which are some of the latest synthetic generators.
@@ 4.1 Influence of content alignment
In Tab. 2 we show the performance achieved with different dataset alignment strategies, as described in Fig. 3. Note that all variants are trained with standard augmentations, including blurring and JPEG compression, as proposed in [46]. From the Table, we can observe that there is a large gain in terms of balanced accuracy ($\simeq$20%) when moving from a text-driven generation (first row) to a solution where real and fake images share the semantic content, both reconstructed and self-conditioned. The proposed solution that uses a diffusion pass demonstrates further improvement on average across all the evaluation metrics. This is highlighted in Fig. 5, where we show the power spectra evaluated by averaging the difference between the real and reconstructed images and the real and self-conditioned images. We observe that self-conditioned generation introduces forensic artifacts even at the lowest frequencies, indicating a detector trained on such images can exploit inconsistencies on a broader range of frequencies.
Figure 6: Robustness analysis in terms of balanced Accuracy carried out on nine generators of the Synthbuster dataset.
@@ 4.2 Effect of content augmentation
Overall, it is evident from Tab. 2 that augmentation plays a critical role in enhancing model generalization and this can be appreciated especially by looking at balanced accuracy and calibration measures. In fact, adding inpainting to reconstruction increases the accuracy from 80.7 to 83.9, while the joint use of self-conditioning and inpainting grants a significant extra gain, reaching 92.2 or even 96.4 with inpainted++. The most significant gains are observed on DALL·E 2, DALL·E 3 and FLUX that, probably, differ the most from Stable 2.1 in terms of architecture and hence require a stronger augmentation strategy to generalize.
@@
DINOv2+reg LP 80.8 68.5 0.58 .141
DINOv2+reg e2e 99.0 95.2 0.14 .040
DINOv2 e2e 98.4 91.1 0.24 .077
SigLIP e2e 95.4 89.9 0.28 .066
Table 3: We compare our solution, DINOv2+reg trained end-to-end, with linear probing (LP) and also consider alternative architectures, basic DINOv2 and SigLIP.
[…]
CLIP/ViT ProGAN [46] 54.7 45.2 4.85 .525
LDM [10] 63.9 48.5 3.39 .487
Ours 75.2 73.9 0.66 .225
RINE ProGAN [46] 66.1 65.0 5.43 .312
LDM [10] 83.0 75.7 0.53 .132
Ours 89.9 83.2 0.39 .089
Table 4: Ablation study on the influence of the training data (ProGAN, Latent Diffusion and our dataset) on methods used for AI-generated image detection: CLIP [37] and RINE [22].
@@
Table 5: Comparison with SoTA methods in terms of balanced Accuracy and balanced NLL across different generators. Note that AntifakePrompt [8] provides only hard binary labels hence calibration measures cannot be computed. Bold underlines the best performance for each column with a margin of 1%.
In Fig. 6, we analyze the impact of our content augmentation, assessing robustness under various operations: JPEG compression, resizing, and blurring. All three proposed variants of augmentation offer a clear advantage, especially when resizing is applied. The joint use of self-conditioning and inpainting results in the most robust approach.
4.3 Influence of training data and architecture
We conduct additional experiments to gain deeper insights into the impact of the chosen architecture and the proposed training data on the same datasets shown in Tab. 2.
@@ 4.3 Influence of training data and architecture
First we compare our adopted model, DINOv2+reg trained end-to-end, with an alternative fine-tuning strategy that involves training only the final linear layer, known as linear probing (LP) that is largely adopted in the literature [33, 11]. From Tab. 3 we can see that this latter solution does not perform well. One possible explanation is that features from last layer capture high-level semantics, while our dataset is built to exploit low-level artifacts that derive from first and intermediate layers [10, 22]. In the same Table we compare DINOv2+reg with the basic DINOv2 [34] architecture and SigLIP [53] and we can observe that DINOv2 with the use of registers achieves the best performance, probably thanks to the fact that it avoids to discard local patch information [13].
Then, in Tab. 4, we consider a CLIP-based model and the architecture of RINE, and vary the training dataset by including two well known datasets largely used in the literature, one based on ProGAN [46] and the other on Latent Diffusion [10]. We note that our training paradigm achieves the best performance over all the metrics with a very large gain (RINE increases the accuracy from 65% to 83.2%).
Figure 7: Average performance in term of AUC and bAcc on four datasets: Synthbuster, GenImage, FakeInversion, SynthWildX.
5 Comparison with the State-of-The-Art
In this Section, we conduct a comparison with SoTA methods on 27 diverse synthetic generation models. To ensure fairness, we include only SoTA methods with publicly available code and/or pre-trained models. The selected methods are listed in Table 6 and are further described in the supplementary material together with additional experiments. For all the experiments now on, ours refers to the detector trained using inpainted++ augmentation.
@@ 5 Comparison with the State-of-The-Art
[…]
Table 6: AI-generated image detection methods used for comparison and whose code is publicly available. We specify source and size of the training dataset, and whether augmentation is applied.
A first experiment is summarized in Tab. 5 with results given in terms of balanced accuracy and NLL. Most of the methods struggle to achieve a good accuracy, especially on more recent generators. Instead, B-Free obtains a uniformly good performance on all generators, irrespective of the image origin, whether they are saved in raw format or downloaded from social networks, outperforming the second best (see last column) by $+20.7\%$ in terms of bAcc. Then, we evaluate again all methods on GenImage (unbiased), FakeInversion [7], and SynthWildX [11]. As these datasets encompass multiple generators, we only report the average performance in Tab. 7. On these additional datasets, most methods provide unsatisfactory results, especially in the most challenging scenario represented by SynthWildX, with images that are shared over the web. The proposed method performs well on all datasets, just a bit worse on FakeInversion. Finally in Fig. 7 we study how AUC compares with balanced accuracy for all the methods over several datasets. We observe that some methods, like NPR and LGrad, present a clear non-uniform behavior, with very good performance on a single dataset and much worse on the others. This seems to suggest that these methods may not be truly detecting forensic artifacts, instead are rather exploiting intrinsic biases within the dataset. Differently, the proposed method presents a uniform performance across all datasets and a small loss between AUC and accuracy.
Figure 8: Results of SoTA detectors on real and fake images that went viral on internet, analyzing multiple web-scraped versions of each image. The performance is in terms of balanced accuracy evaluated from the initial online post (Log scale). Only detectors with an accuracy over 50% are shown.
Analysis on content shared on-line. Distinguishing real from synthetic images on social networks may be especially challenging due to the presence of multiple re-posting that impair image quality over time. A recent study conducted in [21] analyzed the detector behavior on different instances of an image shared online, showing that the performance degrades noticeably in time due to repeated re-posting. To better understand the impact of our augmentation strategies on such images, we collected a total of 1400 real/fake images that went viral on the web, including several versions of the same real or fake image.
Fig. 8 illustrates the accuracy, which is evaluated over a 100-day period from the time of initial publication, with times on a logarithmic scale. We compare our proposal with the best performing SoTA methods. We can notice that the performance drops after only one day, after which most competitors are stuck below 70%, with the exception of DeFake that achieves around 75%. Only the proposed method, which comprises more aggressive augmentation, is able to ensure an average accuracy around 92% even after many days from the first on-line post.
@@ 7 Conclusions
In this paper, we propose a new training paradigm for AI-generated image detection. First of all, we empirically demonstrate the importance of pairing real and fake images by constraining them to have the same semantic content. This helps to better extract common artifacts shared across diverse synthetic generators. Then we find that using aggressive data augmentation, in the form of partial manipulations, further boosts performance both in term of accuracy and of calibration metrics. This is extremely relevant especially when working in realistic scenarios, such as image sharing over social networks. Our findings emphasize that careful dataset curation and proper training strategy can be more impactful compared to developing more complex algorithms. We hope this work will inspire other researchers in the forensic community to pursue a similar direction, fostering advancements in bias-free training strategies.
@@ Acknowledgments.
In this supplementary material, we report more details about our implementation (Sec. A). Moreover, we briefly describe the state of the art methods we compare to (Sec. B), and give more details about the calibration metrics used in the experiments (Sec. C). We also provide additional ablation results in Sec. D and carry out further experiments on generalization (Sec. E) on additional publicly available datasets. Furthermore, we provide more results on the robustness of our approach compared with SoTA methods (Sec. F).
Appendix A Implementation Details
Training strategy. The proposed model leverages the DINOv2+reg [34, 13] $504\times 504$ image embedding network as its backbone, followed by a fully connected layer. The model is trained end-to-end using the binary cross-entropy loss function on an NVIDIA A100 GPU. The training process employs the ADAM optimizer with a learning rate of 1e-6, a weight decay of 1e-6, and a batch size of 24. During training, the balanced accuracy is evaluated on a validation set every 3435 iterations. Early stopping is applied to prevent overfitting: training is completed if the validation balanced accuracy does not improve by at least 0.1% over five consecutive evaluations.
Test strategy. If the test image is less than 504 pixels, padding is applied after patch embedding. Otherwise, we average the logit score over multiple crops to analyze the whole image.
@@ Appendix A Implementation Details
Training Dataset. Here we give more deatils on how we built our dataset. Starting from the MS-COCO training set, consisting of 118K images with 80 categories of objects, we first discarded images with licenses different than Creative Commons. Before editing the images, we extracted the largest central crop, which allows us to retain most of the semantic content of the original image. We discarded images where objects are not present and ended up with a pristine source of 51,517 images. For content augmentation, we replaced the selected object with an object generated from the same category using the COCO segmentation mask, and from a different category using a rectangular box. We took care to not affect too much the realism of the content, so for the “different category” case the object is changed with one from a similar category, that belongs to the same COCO supercategory. In this scenario, the only exception is the category person, which does not have a supercategory and it is therefore replaced with a random object. As mentioned in the main paper, besides the default inpainting, we also consider a version where we take the pixels of the object from the generated image, and the pixels of the background from the original one. We did the same with the self-conditioned image, restoring the background with original pixels. Therefore, we ended up with six fake versions for each real image (Fig. 4).
Appendix B SoTA Methods
Below we provide a brief description of the methods we included in the comparison in Section 5. The training datasets used by these methods are indicated in Table 6.
@@ Appendix B SoTA Methods
D3 [3] 0.33 / .156 0.03 / .016 0.27 / .146 0.22 / .120 0.22 / .138 1.07 / .350 0.11 / .048 0.26 / .101 0.32 / .079 0.28 / .105 0.31 / .126
Ours 0.04 / .008 0.01 / .006 0.12 / .044 0.10 / .037 0.02 / .011 0.40 / .149 0.14 / .044 0.17 / .032 0.25 / .043 0.11 / .028 0.14 / .040
Table 8: Ablation study on the influence of the training data. We compare the proposal with the same architecture DINOv2+reg trained on a publicly available dataset, D3 [3], that includes 4 generators from the Stable Diffusion family. Performance are presented in terms of AUC/Accuracy (top) and ECE/NLL (bottom).
@@
AntifakePrompt [8]. It relies on a visual question-answering (VQA) tool, InstructBLIP. The VQA is used with a fixed question, “Is this photo real?", and fine-tuned to provide accurate responses (“Yes" or “No") using a soft prompt tuning technique. Note that the method provides hard binary predictions hence only accuracy can be computed.
@@
CoDE [3]. CoDE trains a Vision Transformer using a contrastive loss similar to CLIP. However, while CLIP aims to learn features for text-image matching, CoDE aims at obtaining an embedding space where real and fake images are effectively separated. We report results using CoDE in combination with the linear classifier.
Appendix C Calibration Metrics